A survey showed that the ratio of telephones to people in the world was 1 : 8 .of the worlds telephones 0.012% were in Ethiopia . In Ethiopia 0.2 % of the population had telephones . At the time of the survey there were 4000 million people in the world. What was the population of Ethiopia?
Answer by richwmiller(17219)

one in eight of the world had telephones
4000*10^6 world population
500*10^6 in the world have phones
.00012*500*10^6 number of phones in Ethiopia =.06 *10^6
.002*x=.06*10^6
x=.06*10^6/.002=30*10^6=30 million people in Ethiopia.